# A Thermodynamic Theory of Learning I: Irreversible Ensemble Transport and Epistemic Costs

### v1

@@ 3.3 Fokker–Planck Dynamics as a Representative Model
The induced ensemble distribution $q_{s}(\theta)$ evolves according to the Fokker–Planck equation
@@ 3.4 Epistemic Free Energy and Entropy Production
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\,T\,\sigma_{s},\qquad\sigma_{s}\geq 0,$ (7)
where $\sigma_{s}$ denotes the instantaneous entropy production rate. Integrating over $s\in[0,1]$ yields
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1},\qquad\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds.$ (8)
@@ 3.5 Free-Energy Decomposition
The first term represents the change in the ensemble-averaged objective, while the second term captures the loss of entropy along the learning trajectory.
@@ 3.6 Gaussian Noise as a Representative Approximation
Gaussian noise is assumed throughout this work for analytical clarity, as it leads to a Fokker–Planck description with a quadratic entropy production rate and a natural Wasserstein-2 geometry. However, Gaussianity is not essential to the conceptual structure of our results. Empirical studies indicate that stochastic gradient noise is often heavy-tailed, in which case the ensemble dynamics are more appropriately described by Lévy-type or fractional diffusion processes. While such dynamics modify the precise form of the transport cost and the associated geometry, they preserve the fundamental features emphasized here: learning induces irreversible probability transport, entropy-like quantities decrease monotonically, and finite-time ensemble transformations incur unavoidable geometric costs. We therefore view the Fokker–Planck formulation as a representative model rather than a literal description of stochastic gradient dynamics.
@@ 4 Epistemic Speed Limit
Let $\Sigma_{0:1}$ denote the entropy production accumulated along the learning trajectory, as defined in Section 3. Then the following statements hold:
@@
    $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:1}.$
@@
    $\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\;\geq\;W_{2}(q_{0},q_{1})^{2}+T\bigl(H[q_{0}]-H[q_{1}]\bigr).$
Moreover, the bound in (2) is tight: equality is achieved when the ensemble trajectory follows a constant-speed Wasserstein geodesic.
@@
$T\,\Sigma_{0:1}=\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
The final inequality follows directly from the algebraic decomposition of the free energy. ∎
@@
$\Sigma_{0:\mathcal{T}}=\frac{1}{T}\int_{0}^{\mathcal{T}}\!\int q_{\tau}(\theta)\,\|v_{\tau}(\theta)\|^{2}\,d\theta\,d\tau=\frac{1}{\mathcal{T}}\,\Sigma_{0:1}.$
This inequality shows that, for a fixed ensemble transformation, the minimal irreversible cost diverges as the available physical training time $\mathcal{T}$ decreases.
Combining the free-energy dissipation identity $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=T\,\Sigma_{0:\mathcal{T}}$ with the physical-time ESL yields the corresponding bound on free-energy drop:

### v2

@@ 3.3 Fokker–Planck Dynamics as a Representative Model
This dynamics can be written in continuity form as
$\partial_{s}q_{s}+\nabla\cdot(q_{s}v_{s})=0,$ (5)
with the associated probability-flow velocity field
$v_{s}=-\nabla\bigl(\Phi+T\log q_{s}\bigr).$ (6)
Substituting this expression for $v_{s}$ into the continuity equation recovers the Fokker–Planck equation below.
@@ 3.4 Epistemic Free Energy and Entropy Production
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\sigma_{s},\qquad\sigma_{s}\geq 0,$ (9)
where $\sigma_{s}$ denotes the instantaneous entropy production rate. A complete derivation of the dissipation identity, including all regularity assumptions and integration-by-parts steps, is provided in Appendix A. Integrating over $s\in[0,1]$ yields
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:1},\qquad\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds.$ (10)
@@ 3.5 Free-Energy Decomposition
The first term represents the change in the ensemble-averaged objective, while the second term accounts for the change in entropy along the learning trajectory (it is negative when entropy decreases).
@@ 3.6 Gaussian Noise as a Representative Approximation
Gaussian noise is assumed throughout this work for analytical clarity, as it leads to a Fokker–Planck description with a quadratic entropy production rate and a natural Wasserstein-2 geometry. However, Gaussianity is not essential to the conceptual structure of our results. Empirical studies indicate that stochastic gradient noise is often heavy-tailed, in which case the ensemble dynamics are more appropriately described by Lévy-type or fractional diffusion processes. While such dynamics modify the precise form of the transport cost and the associated geometry, they preserve the fundamental features emphasized here: learning induces irreversible probability transport, the free energy decreases monotonically, and finite-time ensemble transformations incur unavoidable geometric costs. We therefore view the Fokker–Planck formulation as a representative model rather than a literal description of stochastic gradient dynamics.
@@ 4 Epistemic Speed Limit
Let $\Sigma_{0:1}:=\int_{0}^{1}\sigma_{s}\,ds$ denote the entropy production accumulated along the learning trajectory, where $\sigma_{s}$ is defined in Section 3. Then the following statements hold:
@@
    $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:1}.$ (14)
@@
    $\mathbb{E}_{q_{0}}[\Phi]-\mathbb{E}_{q_{1}}[\Phi]\;\geq\;W_{2}(q_{0},q_{1})^{2}+T\bigl(H[q_{0}]-H[q_{1}]\bigr).$ (16)
Moreover, the bound in (15) is tight: equality is achieved for a constant-speed Wasserstein geodesic between $q_{0}$ and $q_{1}$, although such a trajectory does not necessarily coincide with the Fokker–Planck dynamics for a fixed objective $\Phi$.
@@
$\Sigma_{0:1}=\int_{0}^{1}\!\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta\,ds\;\geq\;W_{2}(q_{0},q_{1})^{2}.$
The final inequality follows by combining the Epistemic Speed Limit with the algebraic decomposition of the free energy. ∎
@@
$\Sigma_{0:\mathcal{T}}=\int_{0}^{\mathcal{T}}\!\int q_{\tau}(\theta)\,\|v_{\tau}(\theta)\|^{2}\,d\theta\,d\tau=\frac{1}{\mathcal{T}}\,\Sigma_{0:1}.$
This scaling shows that, for a fixed ensemble transformation, the minimal irreversible cost diverges as the available physical training time $\mathcal{T}$ decreases.
Combining this scaling with the free-energy dissipation identity $\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\Sigma_{0:\mathcal{T}}$ yields the corresponding finite-time bound on free-energy reduction:
@@
This inequality should be interpreted as a geometric lower bound determined by the available learning time.
@@ 6 Conclusion
Appendix A Monotonicity of the Epistemic Free Energy
In this appendix, we provide a complete derivation showing that the epistemic free-energy functional decreases monotonically along the Fokker–Planck learning dynamics and that its rate of decrease coincides with the epistemic entropy production rate defined in the main text.
Lemma 1 (Free-energy dissipation along Fokker–Planck dynamics).
Let $q_{s}(\theta)$ be a smooth probability density evolving on $s\in[0,1]$ according to the Fokker–Planck equation
$\partial_{s}q_{s}=\nabla\cdot\!\left(q_{s}\nabla\Phi(\theta)\right)+T\,\Delta q_{s},\qquad T>0,$ (18)
where $\Phi(\theta)$ is a time-independent objective function. Define the epistemic free-energy functional
$\mathcal{F}[q]:=\int q(\theta)\,\Phi(\theta)\,d\theta-T\,H[q],\qquad H[q]:=-\int q(\theta)\log q(\theta)\,d\theta.$ (19)
Assume that $q_{s}$ is strictly positive and decays sufficiently fast at infinity (or satisfies appropriate boundary conditions) so that all integrations by parts are justified. Then the free energy satisfies
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\sigma_{s}\leq 0,$ (20)
where
$\sigma_{s}:=\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta,\qquad v_{s}(\theta):=-\nabla\!\left(\Phi(\theta)+T\log q_{s}(\theta)\right),$ (21)
is the epistemic entropy production rate. Consequently,
$\mathcal{F}[q_{0}]-\mathcal{F}[q_{1}]=\int_{0}^{1}\sigma_{s}\,ds.$ (22)
Proof.
We differentiate the free-energy functional along the trajectory $q_{s}$. Using $\frac{d}{ds}\int q\Phi=\int\Phi\,\partial_{s}q$ and $\frac{d}{ds}\int q\log q=\int(1+\log q)\,\partial_{s}q$, we obtain
$\frac{d}{ds}\mathcal{F}[q_{s}]=\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\partial_{s}q_{s}\,d\theta.$ (23)
Substituting the Fokker–Planck equation (18) yields
$\frac{d}{ds}\mathcal{F}[q_{s}]$ $=\int\bigl(\Phi+T(1+\log q_{s})\bigr)\Bigl[\nabla\cdot(q_{s}\nabla\Phi)+T\,\Delta q_{s}\Bigr]\,d\theta.$ (24)
We treat the two terms separately. By integration by parts and the assumed boundary conditions,
$\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\nabla\cdot(q_{s}\nabla\Phi)\,d\theta$ $=-\int q_{s}\,(\nabla\Phi+T\nabla\log q_{s})\cdot\nabla\Phi\,d\theta,$ (25)
$T\int\bigl(\Phi+T(1+\log q_{s})\bigr)\,\Delta q_{s}\,d\theta$ $=-T\int(\nabla\Phi+T\nabla\log q_{s})\cdot\nabla q_{s}\,d\theta.$ (26)
Using $\nabla q_{s}=q_{s}\nabla\log q_{s}$ and combining the two expressions, we obtain
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\int q_{s}(\theta)\,\|\nabla\Phi(\theta)+T\nabla\log q_{s}(\theta)\|^{2}\,d\theta.$ (27)
Introducing the velocity field $v_{s}(\theta):=-\nabla(\Phi(\theta)+T\log q_{s}(\theta))$, this expression reduces to
$\frac{d}{ds}\mathcal{F}[q_{s}]=-\int q_{s}(\theta)\,\|v_{s}(\theta)\|^{2}\,d\theta=-\sigma_{s},$ (28)
which proves (20). Integrating over $s\in[0,1]$ yields (22). ∎